# Secure and Efficient Access Control Framework for Computer-Use Agents via Context Space

### v3

@@ 7.5. Efficacy of Policy Evolution (RQ5)
As shown in Table 4, CSAgent-RF maintains 0% ASR while improving UUA to 54.23%, demonstrating enhanced utility preservation without compromising security guarantees. Figure 7 shows the performance improvements: CSAgent-RF reduces utility drop to 6.76% and latency overhead to 0.53%. Notably, the reduced latency overhead primarily stems from a decrease in the average number of completion steps for some complex tasks. More precise policies enable the agent to execute tasks more efficiently by eliminating unnecessary validation failures and re-planning cycles. Our analysis reveals three primary categories of improvements in the refined context spaces. First, PEF generates more explicit and fine-grained intent definitions. Second, PEF produces more precise constraint conditions and validation logic. Third, PEF relaxes overly conservative security constraints that unnecessarily block benign operations. As a result, CSAgent-RF achieves more accurate runtime intent extraction and policy retrieval, performs more precise validation, and maintains an effective balance between blocking dangerous operations and permitting legitimate actions.
@@ 9. Related Work
Access control and contextual integrity. Traditional access control mechanisms in OS include discretionary access control (DAC) and mandatory access control (MAC) (sandhu1996authentication). SELinux  (smalley2013seandroid; wright2002selinux) exemplifies MAC principles by defining fine-grained policies that govern process interactions and resource access based on security contexts. The theory of contextual integrity (CI) (nissenbaum2004privacy; nissenbaum2009privacy), proposed in more recent years, has been explored by researchers for its applications in existing operating systems (roesner2012userdrivenac; wijesekera2015androidprem), driving the evolution of permission management and access control toward more user-transparent, secure, and flexible approaches.
@@
Access control for LLM agents. For LLM agents, access control focuses on constraining tool calls through pre-execution validation. Most approaches use contextual information to make authorization decisions, following the CI principle. AirGapAgent (bagdasarian2024airgapagent) employs a separate LLM to directly make decisions about data access permissions. GuardAgent (xiang2024guardagent) employs LLMs to generate guardrail code that validates predefined rules before executing actions. VeriSafe Agent (lee2025verisafe) describes rules through Horn clauses and verifies agent behavior based on logic-based reasoning. ShieldAgent (chen2025shieldagent) enforces explicit policy compliance by constructing action-based probabilistic rule circuits for formal verification. QuadSentinel (yang2025quadsentinelsequentsafetymachinecheckable) employs a multi-agent guard team to enforce safety through machine-checkable sequent logic rules derived from natural language policies. Conseca (tsai2025contextualagent) and Progent (shi2025progent) achieve dynamic security policy generation, enabling more autonomous responses to diverse task scenarios. AgentSentinel (hu2025agentsentinelendtoendrealtimesecurity) extends context extraction to the OS, providing real-time protection for CUAs by combining rule-based and LLM-based auditing. These approaches rely on runtime LLM-based security decision-making, suffering from security or performance limitations. AgentSpec (wang2025agentspeccustomizableruntimeenforcement) employs predefined static policies, but adopts a risk-blocking paradigm that struggles with unknown threats and lacks user intent differentiation, limiting policy flexibility.
@@
Other system-level LLM agent protection. Beyond access control, some other forms of system-level protection have been proposed for agents. The Instruction Hierarchy  (wallace2024instruction) and StruQ  (chen2024struq) implement privilege and type separation for LLM inputs to mitigate interference between different input sources. f-secure LLM system  (wu2024systemleveldefenseindirectprompt), RTBAS  (zhong2025rtbasdefendingllmagents), and SAFEFLOW  (li2025safeflowprincipledprotocoltrustworthy) introduce information flow control into LLM agents to prevent prompt injection and privacy leakage. GoEX  (patil2024goex) implements post-facto validation of agent actions and isolates dangerous operations through sandboxing. IsolateGPT  (wu2025isolategpt) employs isolation for multi-LLM application scenarios to prevent interference between LLMs and different applications. ACE (li2025acesecurityarchitecturellmintegrated) decouples agent planning into two phases to defend against untrusted third-party apps. These approaches and CSAgent can collectively form part of a defense-in-depth strategy (website:googlesecagent), working in conjunction to achieve more comprehensive protection.

### v4

@@ 7.5. Efficacy of Policy Evolution (RQ5)
As shown in Table 4, CSAgent-RF maintains 0% ASR while improving UUA to 54.23%, demonstrating enhanced utility preservation without compromising security guarantees. Figure 7 shows the performance improvements: CSAgent-RF reduces utility drop to 6.76% and latency overhead to 0.53%. Across all benchmarks, the overall utility decrease is reduced to 5.42%, with an average latency overhead of only 1.99%. Notably, the reduced latency overhead primarily stems from a decrease in the average number of completion steps for some complex tasks. More precise policies enable the agent to execute tasks more efficiently by eliminating unnecessary validation failures and re-planning cycles. Our analysis reveals three primary categories of improvements in the refined context spaces. First, PEF generates more explicit and fine-grained intent definitions. Second, PEF produces more precise constraint conditions and validation logic. Third, PEF relaxes overly conservative security constraints that unnecessarily block benign operations. As a result, CSAgent-RF achieves more accurate runtime intent extraction and policy retrieval, performs more precise validation, and maintains an effective balance between blocking dangerous operations and permitting legitimate actions.
@@ 9. Related Work
Access control and contextual integrity. Traditional access control mechanisms in OS include discretionary access control (DAC) and mandatory access control (MAC) (Sandhu and Samarati, 1996). SELinux  (Smalley and Craig, 2013; Wright et al., 2002) exemplifies MAC principles by defining fine-grained policies that govern process interactions and resource access based on security contexts. The theory of contextual integrity (CI) (Nissenbaum, 2004, 2009), proposed in more recent years, has been explored by researchers for its applications in existing operating systems (Roesner et al., 2012; Wijesekera et al., 2015), driving the evolution of permission management and access control toward more user-transparent, secure, and flexible approaches.
@@
Access control for LLM agents. For agents, access control focuses on constraining tool calls through pre-execution validation. Most approaches use contextual information to make authorization decisions, following the CI principle. AirGapAgent (Bagdasarian et al., 2024) employs a separate LLM to directly make decisions about data access permissions. GuardAgent (Xiang et al., 2024) employs LLMs to generate guardrail code that validates predefined rules before executing actions. VeriSafe Agent (Lee et al., 2025) describes rules through Horn clauses and verifies agent behavior based on logic-based reasoning. ShieldAgent (Chen et al., 2025) enforces explicit policy compliance by constructing action-based probabilistic rule circuits for formal verification. QuadSentinel (Yang et al., 2025) employs a multi-agent guard team to enforce safety through machine-checkable sequent logic rules derived from natural language policies. Conseca (Tsai and Bagdasarian, 2025) and Progent (Shi et al., 2025a) achieve dynamic security policy generation, enabling more autonomous responses to diverse task scenarios. AgentSentinel (Hu et al., 2025) extends context extraction to the OS, providing real-time protection for CUAs by combining rule-based and LLM-based auditing. These approaches rely on runtime LLM-based security decision-making, suffering from security or performance limitations. AgentSpec (Wang et al., 2025a) employs predefined static policies, but adopts a risk-blocking paradigm that struggles with unknown threats and lacks user intent differentiation, limiting policy flexibility.
@@
Other system-level LLM agent protection. Beyond access control, some other forms of system-level protection have been proposed for agents. The Instruction Hierarchy  (Wallace et al., 2024) and StruQ  (Chen et al., 2024) implement privilege and type separation for LLM inputs to mitigate interference between different input sources. f-secure LLM system  (Wu et al., 2024), RTBAS  (Zhong et al., 2025), and SAFEFLOW  (Li et al., 2025c) introduce information flow control into LLM agents to prevent prompt injection and privacy leakage. GoEX  (Patil et al., 2024a) implements post-facto validation of agent actions and isolates dangerous operations through sandboxing. IsolateGPT  (Wu et al., 2025) employs isolation for multi-LLM application scenarios to prevent interference between LLMs and different applications. ACE (Li et al., 2025a) decouples agent planning into two phases to defend against untrusted third-party apps. These approaches and CSAgent can collectively form part of a defense-in-depth strategy (Google, 2025c), working in conjunction to achieve more comprehensive protection.